If there is no heat losses the heat released by the condensation of x grams of steam at 100c into water at 100c converts y grams of ice at 0c into water at 100c find the ratio of y to x

Grade:8

1 Answers

Adarsh
6 years ago
according to your question,
x.Lv=y.Lf + y.100.s  where Lv and Lf are latent heat of vapourisation and fusion. and s is specific  =>  x.Lv=y(Lf + 100.s) heat of water
=>y/x = Lv/Lf + 100s, putting the values we get y/x = 3..
hope this will help you .
Thanks and Regards 
Adarsh